OpenStudy (anonymous):

what is the square root of 112 and how did you get that?

OpenStudy (beccaboo333):

Do you know how to square something?

OpenStudy (bakonloverk):

Square root is when a number is multiplied by itself to get a total of a different number For example. \[\sqrt{81}\] 9x9 =81 \[\sqrt{36}\] 6x6=36 So to find the square root of 112, you need to figure out which number multiplied by itself equals 112. 112 is not a perfect square so you can simply use a calculator. The square root would be between 10-11

OpenStudy (bakonloverk):

Something in that is wrong.

OpenStudy (the_fizicx99):

\(\ \sqrt{112} \rightarrow \sqrt{16 * 7} \rightarrow 4\sqrt{7} \)

OpenStudy (the_fizicx99):

Ish?

OpenStudy (the_fizicx99):

\(\ \sqrt{16} = 4 \) 7 is prime >.>

OpenStudy (whpalmer4):

Are we 100% certain that the problem is not to find the square root of 121?

OpenStudy (the_fizicx99):

That would be 11 @whpalmer4, there's a possibility because it's a perfect square

OpenStudy (whpalmer4):

That's why I suggested it...a simple typo away from the problem as posted...